# The conformal Laplacian and positive scalar curvature metrics on manifolds with boundary

### v4

@@ Abstract.
We give examples of spin $4$-manifolds with boundary that do not admit metrics of positive scalar curvature and totally geodesic boundary. These manifolds in fact have the stronger property that the conformal Laplacian with appropriate boundary conditions is never positive. The obstruction to the positivity of the conformal Laplacian is given by a real-valued $\xi$-invariant associated to the APS theorem for the twisted Dirac operator. We use analytic techniques related to the prescribed scalar curvature problem in conformal geometry to move beyond earlier work where the metric is a product near the boundary.
@@ 1. Introduction
There is a well developed theory of determining which closed manifolds $M$ have topological obstructions to admitting positive scalar curvature (psc) metrics (see [22]). The corresponding theory for manifolds with boundary $(M,\partial M)$ is less complete. The main results in this paper are both geometric and analytic extensions of known results for psc metrics which are a product near the boundary. On the geometric side, we weaken the product metric assumption to the totally geodesic boundary case. On the analytic side, we prove results relating psc metrics to the positivity of the conformal Laplacian with certain geometric boundary conditions. In particular, an obstruction to a metric with positive conformal Laplacian and totally geodesic boundary is automatically an obstruction to a psc metric with the same boundary condition. As a result, the main technical work in the paper is in geometric analysis, specifically in conformal geometry on manifolds with boundary.
To state the main results, let $R_{g}$ be the scalar curvature of a Riemannian metric $g$ on $(M,\partial M)$ and let $A_{g}$ be second fundamental form of $\partial M$. $\partial M$ is totally geodesic if $A_{g}\equiv 0.$ Let $\eta_{1}$ be the first eigenvalue of the conformal Laplacian with boundary conditions given in (2). For $M$ a spin $4k$-manifold, let $\xi_{\alpha}(\partial M,g|_{\partial M})$ be the $\mathbb{R}$-valued $\xi$-invariant of $\partial M$ associated to the Dirac operator on $\partial M$, with $\alpha$ a unitary representation of $\pi_{1}(\partial M).$
@@
Let $(M,\partial M,g)$ be a spin manifold with $\eta_{1}>0$ and totally geodesic boundary. Assume that $\alpha$ is a unitary representation of $\pi_{1}(\partial M)$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$ In particular, if $M$ admits a metric $g$ with $R_{g}>0$ and $A_{g}\equiv 0$, then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$
@@
Here $\zeta_{1}$ is the lowest eigenvalue of the conformal Laplacian on $\partial M.$ The first implication follows from Cor. 3.3 and the variational definition of $\eta_{1}$, while the second implication uses [27, Thm. 1.2] and the APS index theorem. Therefore, while the most appealing hypotheses to geometers involve the scalar curvature and the second fundamental form, a conformal analytic assumption like $\eta_{1}>0$ is a better/weaker assumption on the Riemannian metric.
@@
The first implication follows from the solution of the Yamabe conjecture: $\eta_{1}>0$ implies $\tilde{g}$ exists. For the second implication, see [21, Thm. 3.5]. In contrast, the use of conformal analytic techniques seems essential in the manifold with boundary case.
For related work, [6] uses $\eta$-invariants to produce examples of psc metrics $g_{1},g_{2}$ on boundary components of a spin manifold $M$ such that $M$ admits no psc metric which is a product metric $dt^{2}\oplus g_{i}$ near the boundary. These metrics have vanishing second fundamental form, so the work in §§2-3 can be thought of as the geometric analysis needed to move beyond the product metric case. Using a combination of Ricci flow and other dimension-specific techniques, [7] completely characterizes the topology of all $3$-manifolds with boundary admitting metrics with $R_{g}>0$ and non-negative mean curvature. In a different approach, [5, Thm. 19] gives a differential topological obstruction in terms of $K$-area to a manifold with $R_{g}>0$ and non-negative mean curvature, and produces an example of a $3$-manifold admitting no such metric. In the other direction, the existence of constant psc metrics on manifolds with boundary goes back to [9, 10] (with [10] corrected in [24]), with recent work by the third author [25, 26, 28]. [19, Thms. 5.3, 5.5] gives topological conditions under which a totally non-spin $M$ has a psc metric with vanishing or positive mean curvature. These results have the advantage of having boundary assumptions only on the mean curvature, rather than the second fundamental form.
As an outline of the paper, in §2 we give the analytic setup. As a first conformal analytic/geometric result, we prove that $\eta_{1}>0,\zeta_{1}>0$ implies the existence of a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}},R_{\tilde{g}|_{\partial M}}$ positive (Thm. 2.1). In §3, we relate the positivity of $\Box_{g}$ on $M$ to geometric and analytic conditions on $\partial M.$ In particular, we replace the assumption that $\zeta_{1}>0$ in Thm. 2.1 with the assumption that $\partial M$ is totally geodesic (Cor. 3.2). We find a geometric obstruction, namely $R_{\bar{g}}\leq 0$, to filling in an $n$-manifold $(X,\bar{g})$ to an $(n+1)$-manifold $(M,g)$ with $\Box_{g}>0$ and $X$ totally geodesic (Prop. 3.3). We similarly find an analytic obstruction to finding a spin manifold $(M,g)$ with $\Box_{g}>0$ and totally geodesic boundary, namely the presence of nontrivial harmonic spinors on $\partial M$ (Cor. 3.4). The transition from the analytic assumption $\zeta_{1}>0$ in §2 to the geometric assumption $A_{g}=0$ in §3 makes the geometric analysis more difficult: compare the simple boundary argument in Thm. 2.1 with the more difficult proof of Cor. 3.2.
@@
We would like to thank the referee for pointing out an error in an earlier version of the paper.
@@ 2. Positive Scalar Curvature and Positivity of the conformal Laplacians
Throughout this paper, let $(M^{n},g)$ be a connected, compact, orientable smooth Riemannian manifold with smooth boundary $\partial M$, $n=\dim M\geqslant 3$. In this section, we give analytic conditions, namely the positivity of the conformal Laplacians on $M$ and $\partial M$, under which there exists a metric $\tilde{g}\in[g]$, the conformal class of $g$, such that the scalar curvatures $R_{\tilde{g}}$ on $M$ and $R_{\imath^{*}\tilde{g}}$ on $\partial M$ are positive, for $\imath:\partial M\to M$ the inclusion (Thm. 2.1). This result is used in §3 (Cor. 3.2, Cor. 3.4), which in turn are used in §4 (Thm. 4.1).
@@ 2.1. Notation and Basic Setup
other sources define $\Box_{g}$ to be $a^{-1}$ times our $\Box_{g}.$ We use a Robin boundary operator on $\partial M$:
@@
Here ${\rm sgn}(\zeta_{1})={\rm sgn}(\lambda(\partial M,i^{*}g))$ by [18, Lemma 2.5], and both signs are conformal invariants.
@@ 2.2. The sign of $R_{g}$ and $R_{\imath^{*}g}$
For an appropriate choice of $f$, we proved that the metric $\tilde{g}=u^{p-2}g=u^{4/(n-2)}g$ has $R_{\tilde{g}}=R$ on $M$ and prescsribed scalar curvature $R_{\imath^{*}\tilde{g}}$ on $\partial M$.
@@
By [28, Thm. 4.5], if $\eta_{1}>0$, then for any $R\in\mathcal{C}^{\infty}(M)$ and non-negative function $f\in\mathcal{C}^{\infty}(\partial M)$, there exists a positive, smooth function $u$ and a positive constant $c$ such that
(5) $-a\Delta_{g}u+R_{g}u=cRu^{p-1}\;{\rm in}\;M,u=f\;{\rm on}\;\partial M.$
By the solution of the Yamabe problem on closed manifolds, $\zeta_{1}>0$ implies there exists $\phi\in C^{\infty}(\partial M),\phi>0$, such that the metric $g_{1}=\phi^{\frac{n+1}{n-3}}\imath^{*}g$ has $R_{g_{1}}>0$ on $\partial M$, provided $n\geqslant 4$. When $n=3$, there exists a conformal change $\imath^{*}g\mapsto e^{2\varphi}\imath^{*}g$ such that the Gaussian curvature $R_{e^{2\varphi}\imath^{*}g}>0$ on $\partial M$ [18].
Take any smooth function $R>0$ on $M$, and set
(6) $f=\begin{cases}\left(\phi^{\frac{n+1}{n-3}}\right)^{\frac{n-2}{4}},&n\geqslant 4,\\ \left(e^{2\varphi}\right)^{\frac{n-2}{4}},&n=3.\end{cases}$
Using this $R$ and $f$ in (5), we conclude that $\tilde{g}=u^{p-2}g$ has $R_{\tilde{g}}>0$ on $M$ and $R_{\imath^{*}\tilde{g}}>0$ on $\partial M$. ∎
@@ 3. Positivity of the conformal Laplacian: Geometric and Analytic boundary obstructions
A Riemannian manifold with psc metric $g$ and minimal boundary has positive conformal Laplacian: $\eta_{1}>0$. In this section, we conversely assume $\eta_{1}>0$ and discuss geometric implications. For example, in Corollary 3.1 we give a proof that $\eta_{1}>0$ implies the existence of a psc $\tilde{g}\in[g]$ with minimal boundary, a type of conformal converse. If we extend the hypotheses to $\eta_{1}>0$ and $\partial M$ is totally geodesic, then we find $\tilde{g}\in[g]$ with $R_{\imath^{*}\tilde{g}}>0$ and $\partial M$ still totally geodesic (Proposition 3.2), and (a different) $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ a positive constant on $M$ and $R_{\imath^{*}\tilde{g}}$ a positive constant on $\partial M$ (Theorem 3.3).
These results can be stated as boundary geometric obstructions to $\eta_{1}>0$. For example, Proposition 3.3 states that a closed manifold $(X,\bar{g})$ with $R_{\bar{g}}\leq 0$ cannot be the boundary of a manifold $(M,g)$ with $\eta_{1}>0$ and $\imath^{*}g$ conformal to $\bar{g}.$ We also obtain a boundary analytic obstruction in Corollary 3.4: if a spin manifold has $\eta_{1}>0$ and totally geodesic boundary, then the boundary has no harmonic spinors.
The proofs of these results are much simpler on the class of metrics which are products near the boundary. Thus this section highlights the work necessary to pass from the product metric case to the totally geodesic case.
@@
$R_{g}=2\text{Ric}_{g}(\nu_{g},\nu_{g})+R_{\imath^{*}g}+\lvert A_{g}\rvert^{2}-h_{g}^{2}\;{\rm on}\;\partial M.$
@@
The analysis of this equation under conformal change is key to the results in this section.
3.1. Results on Ric and $A_{g}$
We want to analyze (7) under the assumption that $R_{g}>0.$ This is always possible if $\eta_{1}>0$ and the boundary is minimal or totally geodesic (Thm. 3.1).
@@ 3.1. Results on Ric and $A_{g}$
Assuming $\eta_{1}>0$ and good boundary geometry, we show it is possible to alter a metric within its conformal class to have positive scalar curvature.
@@
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally geodesic, resp. minimal with $n\geqslant 3$, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $\partial M$ remains totally geodesic, resp. minimal, for $\tilde{g}$.
@@
We treat the case where $\partial M$ is totally geodesic; the minimal boundary case is similar. By [9, §1], the eigenvalue problem for the conformal Laplacian
@@
Again by [9, §1], the eigenvalue problem for the conformal Laplacian
@@
We now study (7) under conformal change as in (9). Set $\text{Ric}_{g,nn}=\text{Ric}_{g}(\nu_{g},\nu_{g})$, and similarly for $\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}.$ In terms of $\phi$, the well-known transformation laws for the Ricci curvature tensor and the scalar curvature are

### v5

@@ Abstract.
We give examples of spin $4$-manifolds with boundary that do not admit metrics of positive scalar curvature and minimal or even mean convex boundary. These manifolds in fact have the stronger property that the conformal Laplacian with appropriate boundary conditions is never positive. The obstruction to the positivity of the conformal Laplacian is given by a real-valued $\xi$-invariant associated to the APS theorem for the twisted Dirac operator. We use analytic techniques related to the prescribed scalar curvature problem in conformal geometry to move beyond earlier work where the metric is a product near the boundary.
@@ 1. Introduction
There is a well developed theory of determining which closed manifolds $M$ have topological obstructions to admitting positive scalar curvature (psc) metrics (see [27]). The corresponding theory for manifolds with boundary $(M,\partial M)$ is less complete. The main results in this paper are both geometric and analytic extensions of known results for psc metrics which are a product near the boundary. On the geometric side, we weaken the product metric assumption to the totally umbilic boundary case. On the analytic side, we prove results relating psc metrics to the positivity of the conformal Laplacian with certain geometric boundary conditions. The geometric and analytic results are related: for example, an obstruction to a metric with positive conformal Laplacian and minimal boundary is automatically an obstruction to a psc metric with the same boundary condition. As a result, the main technical work in the paper is in geometric analysis, specifically in conformal geometry on manifolds with boundary.
To state the main results, let $R_{g}$ be the scalar curvature of a Riemannian metric $g$ on $(M,\partial M)$ and let $A_{g}$ be second fundamental form of $\partial M$. $\partial M$ is totally geodesic if $A_{g}\equiv 0$, totally umbilic if $A_{g}=\lambda g$ for some $\lambda\in C^{\infty}(\partial M),$ minimal if ${\rm Tr}(A_{g})\equiv 0$, and mean convex if ${\rm Tr}(A_{g})$ does not point inward anywhere on the boundary. Let $\eta_{1}$ be the first eigenvalue of the conformal Laplacian $\Box_{g}$ with boundary conditions given in (2). For $M$ a spin $4k$-manifold, let $\xi_{\alpha}(\partial M,g|_{\partial M})$ be the $\mathbb{R}$-valued $\xi$-invariant of $\partial M$ associated to the Dirac operator on $\partial M$, with $\alpha$ a unitary representation of $\pi_{1}(\partial M).$
@@
Let $(M,\partial M,g)$ be a spin manifold with $\eta_{1}>0$ and totally umbilic boundary. Assume that $\alpha$ is a unitary representation of $\pi_{1}(\partial M)$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$ In particular, if $M$ admits a metric $g$ with $R_{g}>0$ and $A_{g}\equiv 0$, then $\xi_{\alpha}(\partial M,g|_{\partial M})=0.$
@@
Here $\zeta_{1}$ is the lowest eigenvalue of the conformal Laplacian on $\partial M.$ The first implication follows from the variational definitions of $\eta_{1}$ and $\zeta_{1}$, while the second implication uses [32, Thm. 1.2] and the APS index theorem. Therefore, while the most appealing hypotheses to geometers involve the scalar curvature and the second fundamental form, a conformal analytic assumption like $\eta_{1}>0$ is a better/weaker assumption on the Riemannian metric.
@@
The first implication follows from the solution of the Yamabe conjecture: $\eta_{1}>0$ implies $\tilde{g}$ exists with $R_{\tilde{g}}$ positive. For the second implication, see [26, Thm. 3.5]. In contrast, the use of conformal analytic techniques seems essential in the manifold with boundary case.
For related work, [8] uses $\xi$-invariants to produce examples of psc metrics $g_{1},g_{2}$ on boundary components of a spin manifold $M$ such that $M$ admits no psc metric which is a product metric $dt^{2}\oplus g_{i}$ near the boundary. [2] gives conditions under which a psc metric has a deformation to a psc metric which is a product near the boundary, and [1] discusses when a psc metric on a closed manifold $X$ can be filled in to a psc metric on a manifold $W$ with $\partial W=X$. [11] uses $\xi$-invariants as an obstruction to the filling problem in conformally flat (e.g., hyperbolic) geometry. The metrics in [1, 2, 8] have vanishing second fundamental form and hence totally umbilic and minimal boundary, so the work in §3 can be thought of as the geometric analysis needed to move beyond the product metric case. Using a combination of Ricci flow and other dimension-specific techniques, [9] completely characterizes the topology of all $3$-manifolds with boundary admitting metrics with $R_{g}>0$ and non-negative mean curvature. In a different approach, [7, Thm. 19] gives a differential topological obstruction in terms of $K$-area to a manifold with $R_{g}>0$ and non-negative mean curvature, and produces an example of a $3$-manifold admitting no such metric. In the other direction, the existence of constant psc metrics on manifolds with boundary goes back to [13, 14] (with [14] corrected in [29]), with recent work by the third author [30, 31, 33]. [24, Thms. 5.3, 5.5] gives topological conditions under which a totally non-spin $M$ has a psc metric with vanishing or positive mean curvature. These results have the advantage of having boundary assumptions only on the mean curvature, rather than the second fundamental form.
As an outline of the paper, in §2 we give the analytic setup. As a first conformal analytic/geometric result, we prove that $\eta_{1}>0,\zeta_{1}>0$ implies the existence of a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}},R_{\tilde{g}|_{\partial M}}$ positive (Thm. 2.1). In §3, we relate the positivity of $\Box_{g}$ on $M$ to geometric and analytic conditions on $\partial M.$ In particular, we replace the condition $\zeta_{1}>0$ in Thm. 2.1 with the assumption that $\partial M$ is totally umbilic (Thm. 3.1). We find a geometric obstruction, namely $R_{\bar{g}}\leq 0$, to filling in an $n$-manifold $(X,\bar{g})$ to an $(n+1)$-manifold $(M,g)$ with $\Box_{g}>0$ and $X$ totally umbilic (Cor. 3.2). We similarly find an analytic obstruction to finding a spin manifold $(M,g)$ with $\Box_{g}>0$ and totally umbilic boundary, namely the presence of nontrivial harmonic spinors on $\partial M$ (Cor. 3.4). The transition from the analytic assumption $\zeta_{1}>0$ in §2 to the geometric assumption (totally umbilic boundary) in §3 makes the geometric analysis more difficult: compare the simple boundary argument in Thm. 2.1 with the more difficult proof of Thm. 3.1.
@@
We would like to thank the referee both for pointing out a serious error in an earlier version of the paper and for encouraging us to replace the totally geodesic boundary condition with the weaker totally umbilic condition in §§3,4.
@@ 2. Positive Scalar Curvature and Positivity of the conformal Laplacians
Throughout this paper, let $(M^{n},g)$ be a connected, compact, orientable smooth Riemannian manifold with smooth boundary $\partial M$, $n=\dim M\geqslant 3$. In this section, we give analytic conditions, namely the positivity of the conformal Laplacians on $M$ and $\partial M$, under which there exists a metric $\tilde{g}\in[g]$, the conformal class of $g$, such that the scalar curvatures $R_{\tilde{g}}$ on $M$ and $R_{\imath^{*}\tilde{g}}$ on $\partial M$ are positive, for $\imath:\partial M\to M$ the inclusion (Thm. 2.1). In §3, we will replace the analytic boundary condition with geometric boundary conditions (Prop. 3.1, Thm. 3.1), which is then used in §4 (Thm. 4.1).
@@ 2.1. Notation and Basic Setup
other sources define $\Box_{g}$ to be $a^{-1}$ times our $\Box_{g}.$ We use a Robin boundary operator on $\partial M$ as in [13]:
@@
Here ${\rm sgn}(\zeta_{1})={\rm sgn}(\lambda(\partial M,i^{*}g))$ by [23, Lemma 2.5], and both signs are conformal invariants.
@@ 2.2. The sign of $R_{g}$ and $R_{\imath^{*}g}$
For an appropriate choice of $f$, we proved that the metric $\tilde{g}=u^{p-2}g=u^{4/(n-2)}g$ has $R_{\tilde{g}}=R$ on $M$ and prescribed scalar curvature $R_{\imath^{*}\tilde{g}}$ on $\partial M$.
@@
Let $\tilde{\eta}_{1}$ be the first eigenvalue of the conformal Laplacian $\Box_{g}$ with vanishing Dirichlet boundary condition, i.e., there exists $\varphi\in C^{\infty}(M)$ with $\Box_{g}\varphi=\tilde{\eta}_{1}\varphi\;{\rm in}\;M,\varphi=0\;{\rm on}\;\partial M.$ By [33, Prop. 4.3], $\tilde{\eta}_{1}\geqslant\eta_{1}>0$. Given $f\in\mathcal{C}^{\infty}(\partial M)$, $f>0$, there exists a positive, smooth function $u$ and a positive constant $\lambda$ such that
(5) $-a\Delta_{g}u+R_{g}u=\lambda u^{p-1}\;{\rm in}\;M,u=f\;{\rm on}\;\partial M,$
by [33, Thm. 4.5].
By the solution of the Yamabe problem on closed manifolds, $\zeta_{1}>0$ implies there exists $\phi\in C^{\infty}(\partial M),\phi>0$, such that the metric $g_{1}=\phi^{\frac{4}{n-3}}\imath^{*}g$ has $R_{g_{1}}>0$ on $\partial M$, provided $n\geqslant 4$. When $n=3$, there exists a conformal change $g_{1}=e^{2\varphi}\imath^{*}g$ with Gaussian curvature $R_{g_{1}}>0$ on $\partial M$ [23].
Set
(6) $f=\begin{cases}\left(\phi^{\frac{4}{n-3}}\right)^{\frac{n-2}{4}},&n\geqslant 4,\\ \left(e^{2\varphi}\right)^{\frac{n-2}{4}},&n=3.\end{cases}$
Using this $f$ in (5), we conclude that $\tilde{g}=u^{p-2}g$ has $R_{\tilde{g}}>0$ on $M$ and $R_{\imath^{*}\tilde{g}}>0$ on $\partial M$. ∎
@@ 3. Positivity of the conformal Laplacian: Geometric and Analytic boundary obstructions
A Riemannian manifold with psc metric $g$ and minimal boundary has positive conformal Laplacian: $\eta_{1}>0$. In this section, we conversely assume $\eta_{1}>0$ and discuss geometric implications. For example, in Cor. 3.1 we give a proof that $\eta_{1}>0$ implies the existence of a psc $\tilde{g}\in[g]$ with minimal boundary, a type of conformal converse. If we extend the hypotheses to $\eta_{1}>0$ and $\partial M$ is totally geodesic, then we find $\tilde{g}\in[g]$ with $R_{\imath^{*}\tilde{g}}>0$ and $\partial M$ still totally geodesic (Prop. 3.1). With more work, we can weaken the boundary condition to totally umbilic and find $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ and $R_{\imath^{*}\tilde{g}}$ positive (Thm. 3.1).
These results can be stated as boundary geometric obstructions to $\eta_{1}>0$. For example, Cor. 3.2 states that a closed manifold $(X,\bar{g})$ with $R_{\bar{g}}\leq 0$ cannot be the boundary of a manifold $(M,g)$ with $\eta_{1}>0$ and $\imath^{*}g$ conformal to $\bar{g}.$ We also obtain a boundary analytic obstruction in Cor. 3.4: if a spin manifold has $\eta_{1}>0$ and totally geodesic boundary, then the boundary has no harmonic spinors.
The proofs of these results are much simpler on the class of metrics which are products near the boundary. Thus this section highlights the work necessary to pass first from the product metric case to the totally geodesic case and then to the totally umbilic case.
@@
(7) $R_{g}=2\text{Ric}_{g}(\nu_{g},\nu_{g})+R_{\imath^{*}g}+\lvert A_{g}\rvert^{2}-h_{g}^{2}\;{\rm on}\;\partial M.$
@@
The analysis of this equation under conformal change is key to the results in this section. With extra work, we can weaken the boundary condition to totally umbilic; see §3.2.
3.1. Preliminary results on positivity with totally geodesic or minimal boundary
We collect some known or folklore results about positivity of scalar curvature if $\eta_{1}>0$ and the boundary is minimal or totally geodesic. Proofs are included for completeness.
@@ 3.1. Preliminary results on positivity with totally geodesic or minimal boundary
Assuming $\eta_{1}>0$ and good boundary geometry, we can alter a metric within its conformal class to have positive scalar curvature.
@@
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally geodesic, resp. minimal with $n\geqslant 3$, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $\partial M$ remains totally geodesic, resp. minimal, for $\tilde{g}$. If $\partial M$ is totally geodesic, then $R_{i^{*}\tilde{g}}>0.$
@@
We treat the case where $\partial M$ is totally geodesic; the minimal boundary case is similar. By [13, §1], the eigenvalue problem for the conformal Laplacian
@@
For the last statement, it is standard that a collar of $\partial M$ has coordinates $(x,t)\in\partial M\times[0,1)$ such that $\tilde{g}(x,t)=h(x,t)\oplus dt^{2}$, for some family of metrics $h(x,t)$ on $\partial M$. If $\partial M$ is totally geodesic, a straightforward calculation in these coordinates gives $R_{\tilde{g}}|_{\partial M}=R_{i^{*}\tilde{g}}.$
@@
Again by [13, §1], the eigenvalue problem for the conformal Laplacian
@@
3.2. Positivity with totally umbilic boundary
In this subsection, we weaken the totally geodesic boundary assumption in Prop. 3.1 to totally umbilic (Thm. 3.1). The main point is that we can still guarantee that $R_{i^{*}\tilde{g}}>0$ on $\partial M.$
Definition 3.1.
Let $(M^{n},\partial M,g)$ be a compact Riemannian manifold with boundary. $\partial M$ is totally umbilic if the second fundamental form $A_{g}$ is proportional to the first fundamental form $g$, i.e.
(13) $A_{g}(X,Y)=\lambda_{p}g(X,Y),\forall X,Y\in T_{p}\partial M,p\in\partial M.$
Considering $A_{g}$ as a linear transformation on $T_{p}\partial M$, $\partial M$ is totally umbilic iff all eigenvalues of $A_{g}$ are equal to $\lambda_{p}$. The mean curvature is then
(14) $h_{g}(p)=(n-1)\lambda_{p}.$
From this, it is immediate that a boundary is totally geodesic iff it is minimal and totally umbilic. The Clifford torus is a minimal submanifold of $S^{3}$ which is not totally umbilic, and shows that the positivity of scalar curvature on a manifold with minimal boundary does not imply positivity of the scalar curvature on the boundary. [10, Ex. 4.4] gives an example of a totally umbilic submanifold which is not minimial.
Being totally umbilic is a conformally invariant condition.
Proposition 3.2.
[11, Prop. 2.1] The multiplicities of the eigenvalues of $A_{g}$ are conformal invariants. In particular, $g$ is totally umbilic iff $\tilde{g}\in[g]$ is totally umbilic.
We first relate $R_{g}$ to $R_{i*g}$ in the totally umbilic case. Choose an othonormal local frame $e_{1},\dotso,e_{n}=\nu_{g}$ at a fixed point $p\in\partial M.$ Applying (13 and (14) at $p$, we have
$h_{g}^{2}(p)-\lvert A_{g}\rvert(p)^{2}=\left((n-1)\lambda_{p}\right)^{2}-\sum_{i,j=1}^{n-1}\left(\lambda_{p}g(e_{i},e_{j})\right)^{2}=(n-1)(n-2)\lambda_{p}^{2}\geqslant 0.$
By Gauss-Codazzi (7),
(15) $R_{\imath^{*}g}(p)=R_{g}(p)-2\text{Ric}_{g}(\nu_{g},\nu_{g})(p)+h_{g}^{2}(p)-\lvert A_{g}\rvert(p)^{2}\geqslant R_{g}(p)-2\text{Ric}_{g}(\nu_{g},\nu_{g})(p).$
In particular, $R_{g}-2\text{Ric}_{g}(\nu_{g},\nu_{g})>0$ on $\partial M$ implies $R_{\imath^{*}g}>0$ on $\partial M$.
We therefore focus on the sign of $R_{\tilde{g}}-2\text{Ric}_{\tilde{g}}(\nu_{\tilde{g}},\nu_{\tilde{g}})$ on $\partial M$ for some metric $\tilde{g}\in[g]$. Set $\text{Ric}_{g,nn}=\text{Ric}_{g}(\nu_{g},\nu_{g})$, and similarly for $\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}.$ In terms of $\phi$, the well-known transformation laws for the Ricci curvature tensor and the scalar curvature are